OpenStudy (anonymous):

Factor the perfect square trinomial completely: 16y^2+56y+49

OpenStudy (anonymous):

(4y)² + 2(4y)(7) + (7)² using the identity a² + 2ab + b² = (a + b)² (4y)² + 2(4y)(7) + (7)² = (4y + 7)² so 16y^2+56y+49 = (4y + 7)(4y + 7)